Sum of two numbers is 95. If one exceeds the other by 15, find the numbers.

 

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear student

Let first number be x

So, second number will be x+15

As per the given question;
x+x+15=95
x+x=95-15
2x=80
x=80/2
x=40

x=40

First number = 40

Second number =x+15=40+15=55

Hence, the numbers are 40 and 55

Thanks